Polynomial Inequalities:

10) 6 + 5x - 6x^ 2 > 0

show number line and answer in interval notation


1
Expert's answer
2021-09-24T14:50:39-0400

"6x^2-5x-6=0"


"x=\frac{5\pm \sqrt{25+144}}{12}"


"x_1=-2/3,\ x_2=3/2"


"6x^2-5x-6=(3x+2)(2x-3)"


"(3x+2)(2x-3)<0"


"x\isin (-2/3,3/2)"